# BeamformNet: Deep Learning-Based Beamforming Method for DoA Estimation via Implicit Spatial Signal Focusing and Noise Suppression

### v1

@@ Abstract
Deep learning–based direction-of-arrival (DoA) estimation has gained increasing popularity. A popular family of DoA estimation algorithms is beamforming methods, which operate by constructing a spatial filter that is applied to array signals. However, these spatial filters obtained by traditional model-driven beamforming algorithms fail under demanding conditions such as coherent sources and a small number of snapshots. In order to obtain a robust spatial filter, this paper proposes BeamformNet—a novel deep learning framework grounded in beamforming principles. Based on the concept of optimal spatial filters, BeamformNet leverages neural networks to approximately obtain the optimal spatial filter via implicit spatial signal focusing and noise suppression, which is then applied to received signals for spatial focusing and noise suppression, thereby enabling accurate DoA estimation. Experimental results on both simulated and real-world speech acoustic source localization data demonstrate that BeamformNet achieves state-of-the-art DoA estimation performance and has better robustness.
@@ I Introduction
Direction of arrival (DoA) estimation, a fundamental task in array signal processing, aims to locate sources using signals received by an array of sensors [benesty2017fundamentals]. Classical model-driven methods achieve strong performance under ideal conditions [barabell1983improving, bienvenu1983optimality, schmidt1986multiple, roy2002esprit, capon2005high, malioutov2005sparse]. However, in practice, these methods suffer from performance degradation due to signal coherence, array mismatch, limited snapshots, and so on [krim1996two, tuncer2009classical]. Recently, deep learning has advanced DoA estimation owing to its data-driven nature [papageorgiou2021deep, lee2022deep, merkofer2022deep, gao2023gridless, merkofer2023music, shmuel2023deep, chen2024sdoa, shmuel2024subspacenet, xu2025deep, gast2025dcd]. Despite their black-box nature, high complexity, and limited interpretability, neural networks effectively handle non-ideal conditions where traditional model-driven methods struggle, demonstrating greater robustness in demanding scenarios.
@@
Beamforming methods are a popular family of traditional model-driven DoA estimation algorithms, including conventional beamforming (CBF) [krim1996two], MVDR [capon2005high] and so on. At its core, beamforming constructs a spatial filter applied to array signals to estimate DoA [van1988beamforming]. Based on the idea of spatial filtering and the sparsity of the signal in space, We defined an optimal spatial filter $\bm{B^{*}}$ that focuses on the target signal while suppressing noise and interference, enabling accurate DoA estimation. under the assumptions of narrowband far-field signals and fewer sources than array elements, we show the existence of an optimal spatial filter.
@@
Inspired by deep learning, traditional beamforming, and the optimal spatial filter, we propose a deep learning-based beamforming method to approximate the optimal spatial filter $\bm{B^{*}}$ and apply it to received array signals for target focusing and noise suppression, enabling accurate DoA estimation. Under the established existence of $\bm{B^{*}}$, the universal approximation capability of deep networks [cybenko1989approximation, hornik1989multilayer, leshno1993multilayer, kidger2020universal, yarotsky2022universal] and extensive validation experiments, our method is theoretically grounded and experimentally validated, providing an interpretable foundation and clarifying the theoretical upper bound. Our main contributions are as follows:
@@
  • Open Source. Code is publicly available on GitHub to foster community advancement  [^1] https://github.com/colaudiolab/BeamformNet.
@@ II-A Deep Learning-Based DoA Estimation
Figure 1: The overall structure of BeamformNet.
@@ II-B Beamforming
Beamforming-based DoA estimation methods fall into static [krim1996two, capon2005high] and dynamic [zheng2023interpretable, deng2025spatial] categories: the former uses fixed closed-form expression filters, while the latter employs iterative optimization. They can also be classified as non-adaptive [krim1996two] or adaptive [capon2005high, deng2025spatial], depending on whether filter weights are adjusted based on received signals. In terms of implementation methodology, existing approaches can be divided into model-driven methods [krim1996two, capon2005high, deng2025spatial] and neural network-based methods [al2022review, zheng2023interpretable, hamza2024sparse]. Our method is a dynamic, adaptive beamforming approach implemented via deep neural networks: its dynamic nature arises from iterative data training, and its adaptivity comes from generating spatial filters in response to input array signals.
@@ III-C Beamforming and the Existence of the Optimal Spatial Filter
DoA is estimated from the spectral peak of the spatial energy spectrum $\bm{P}=[p_{1},...,p_{i},..,p_{R}]^{T}$, where $p_{i}=\bm{{b_{i}}^{H}XX^{H}b_{i}}$ with $p_{i}$ is the $i$-th element of $\bm{P}$ and $\bm{b_{i}}$ is the $i$-th column vector of matrix $\bm{B}^{H}$. For example, in the MVDR algorithm,
@@ IV-A Context Extraction in Time and Spatial Domains
To preserve the phase, BeamformNet splits $\bm{A}$ and $\bm{X}$ into real and imaginary parts, concatenates them, and feeds each into separate bidirectional RNNs to extract contextual features in the time and spatial domains in order to adaptively generate the spatial filter $\bm{B}$. Some works [zhou2018direction, su2024co, ma2024doa] show that virtual elements can enhance element dimensions for better DoA estimation. Accordingly, we expand the element dimensions of $\bm{A}$ and $\bm{X}$ from $2M$ to $E$ ($E>2M$) after passing through the RNN, yielding an enriched representation that captures both temporal and spatial contexts with additional element dimensions:
@@ IV-C Spatial Filter
After obtaining the spatial energy spectrum $\bm{P}$, one possible approach is to directly learn according to Formula (12) and calculate the error between the obtained spatial energy spectrum $\bm{P}$ and the sparse spatial signal $\bm{S}$, such as using MSE. Because our goal is to solve DoA, we are not interested in energy. So we directly solve for the direction in which the target angle is located. The DoAs obtained by searching the spatial energy spectrum $\bm{P}$ through spectral peaks cannot be directly used to compute gradients for backpropagation. In order to enable BeamformNet to backpropagate gradients for training, we apply the tanhtanh activation function to $\bm{P}$. Since each element in $\bm{P}$ lies in the range $0,+\infty)$, applying tanhtanh maps the signal energy in each grid cell to a probability distribution over $[0,1]$. Accordingly, BeamformNet formulates model training as a multi-label classification problem, which maximizes the probability of the grid where the target angle is located.
The above process uses the tanhtanh function (monotonically increasing) to convert $\bm{P}$ into probability and utilizes the sparsity of $\bm{S}$ for classification, which implies that the direction energy of the signal in $\bm{P}$ is higher (signal focusing), while other directions are lower or zero (noise suppression). It is equivalent to letting the network implicitly learn Formulas (13) and (14) (see Section VI-I for more details), although our network does not explicitly learn these two formulas.
@@ VI-C Signal-to-Noise Ratio
The SNR measures signal strength relative to noise. A higher SNR indicates better signal quality and more accurate information transmission. We studied SNR’s impact on BeamformNet under $M=16$, $K=4$, $T=200$, and incoherent sources, comparing it with other methods. Figure 7 shows that DA-MUSIC performs best at low SNRs (-20 dB and -15 dB), while SubspaceNet and BeamformNet show comparable results. When the SNR exceeds -15 dB, BeamformNet outperforms all other methods, especially at high SNRs. We attribute BeamformNet’s weaker low-SNR performance to its direct interaction with the noisy signals $\bm{X}$ ($\bm{P}=\bm{BX}$) without any denoising. Adding a denoising module to improve BeamformNet robustness in low-SNR scenarios represents a promising area for future research. Nevertheless, overall, BeamformNet displays strong performance across the range of SNR conditions.
@@ VII Conclusion
In this paper, we propose a novel method, BeamformNet, for DoA estimation. The BeamformNet architecture is inspired by model-driven beamforming algorithms, and the learning framework of BeamformNet is theoretically grounded to approximately achieve optimal DoA estimation by approximating the optimal spatial filter. Extensive experimental results demonstrate that BeamformNet shows robustness and outperforms existing methods across a wide range of scenarios—both in simulated environments and in challenging real-world acoustic conditions involving near-field broadband speech signals. These findings validate the effectiveness of approximating the optimal solution through deep neural network-based learning. Future work will focus on further improving the model’s performance under extreme conditions, such as single-snapshot and very low-SNR scenarios.
@@
[Uncaptioned image] Kele Xu (Senior Member, IEEE) received the doctorate degree from Paris VI University, Paris, France, in 2017. He is currently an Associate Professor with the School of Computer Science, National University of Defense Technology, Changsha, China. His research interests include audio signal processing, machine learning, and intelligent software systems. He is the associate editor for IEEE Transactions on Circuits and Systems for Video Technology and Guest editor for Science Partner Journal Cyborg and Bionic Systems. He has (co-)authored more than 100 publications in peer reviewed journals and conference proceedings, including TASLP, TMI, ICML, CVPR, NeuRIPS, ICLR, AAAI, IJCAI, ASE, ACM MM, ICASSP.

### v2

@@ Abstract
Direction-of-arrival (DoA) estimation is a cornerstone of array signal processing, with beamforming methods representing a prominent family of techniques that construct spatial filters to enhance target signals while suppressing noise and interference. Traditional model-driven beamformers, such as Conventional Beamforming (CBF) and Minimum Variance Distortionless Response (MVDR), excel under ideal conditions but degrade in challenging scenarios involving coherent sources, limited snapshots, array mismatches, or low signal-to-noise ratios (SNRs). To address these limitations, we introduce BeamformNet, a novel deep learning framework that leverages recurrent neural networks (RNNs) and attention mechanisms to approximate an optimal spatial filter, grounded in beamforming principles and signal sparsity in the spatial domain. BeamformNet implicitly performs spatial signal focusing and noise suppression by processing the array manifold and received signals through bidirectional RNNs for contextual feature extraction, followed by an attention module for adaptive fusion and a projection layer to generate the spatial filter. This filter is applied to the received signals to yield a sparse spatial energy spectrum, enabling accurate DoA estimation via peak search after thresholding. Theoretically, we prove the existence of the optimal spatial filter under assumptions of narrowband far-field signals and fewer sources than array elements, with neural networks approximating it via the universal approximation theorem. The framework formulates the optimal spatial filter approximation as a multi-label classification task, trained using asymmetric loss to handle class imbalance. Extensive experiments on simulated datasets and real-world benchmarks demonstrate that BeamformNet achieves state-of-the-art performance, outperforming methods like DA-MUSIC, DeepSSE, SubspaceNet, IQResNet, LowSNRNet, and classical MUSIC across varying numbers of sources (K), snapshots (T), SNRs, angular separations, array elements (M), and mismatches. It exhibits superior generalization and robustness, particularly under coherent sources and limited snapshots. BeamformNet’s parameter count scales modestly with M, ensuring efficiency. Code is open-sourced on GitHub to advance the community  [^1] https://github.com/colaudiolab/BeamformNet.
@@ I Introduction
Direction of arrival (DoA) estimation, a fundamental task in array signal processing, aims to locate sources using signals received by an array of sensors [benesty2017fundamentals]. Classical model-driven methods achieve strong performance under ideal conditions [barabell1983improving, bienvenu1983optimality, schmidt1986multiple, roy2002esprit, capon2005high, malioutov2005sparse]. However, in practice, these methods suffer from performance degradation due to signal coherence, array mismatch, limited snapshots, and so on [krim1996two, tuncer2009classical]. Recently, deep learning has advanced DoA estimation owing to its data-driven nature [papageorgiou2021deep, lee2022deep, merkofer2022deep, gao2023gridless, merkofer2023music, shmuel2023deep, chen2024sdoa, shmuel2024subspacenet, xu2025deep, gast2025dcd]. Despite their black-box nature, high complexity, and limited interpretability, neural networks effectively handle non-ideal conditions where traditional model-driven methods struggle, demonstrating greater robustness in demanding scenarios. How to combine interpretability and robustness in complex scenarios has become a key point in the field of neural network-based DoA estimation.
Beamforming methods are a popular family of traditional model-driven DoA estimation algorithms that are limited by model assumptions and are difficult to adapt to complex scenarios, including Conventional BeamForming (CBF) [krim1996two], Minimum Variance Distortionless Response (MVDR) [capon2005high] and so on. At its core, beamforming constructs a spatial filter applied to array signals to estimate DoA [van1988beamforming]. Different beamforming algorithms are essentially derived from distinct optimization objectives. For instance, MVDR beamformer aims to minimize the total output power while maintaining a unity gain in the direction of interest, thereby suppressing interference and noise [capon2005high]. Other approaches, such as Linearly Constrained Minimum Variance (LCMV) [1450747], extend this by imposing multiple linear constraints to shape the beam pattern.
Furthermore, the concept of signal sparsity plays a pivotal role in modern array processing [1614066, 1580791, 1468495, yang2018sparse]. In the spatial domain, sources are typically sparse—meaning the number of active sources is significantly smaller than the number of possible angular grids. This sparsity allows for the formulation of DoA estimation as a sparse signal recovery problem, where the goal is to reconstruct a sparse source vector from compressed array measurements.
@@
Based on the idea of spatial filtering and the sparsity of the signal in space, We defined an optimal spatial filter $\bm{B^{*}}$ that focuses on the target signal while suppressing noise and interference, enabling accurate DoA estimation. under the assumptions of narrowband far-field signals and fewer sources than array elements, we show the existence of an optimal spatial filter. Our optimization objective is to find the optimal spatial filter.
@@
Inspired by deep learning, traditional beamforming, and the optimal spatial filter, we propose a deep learning-based beamforming method to approximate the optimal spatial filter $\bm{B^{*}}$ and apply it to received array signals for target focusing and noise suppression, enabling accurate and robust DoA estimation with interpretability. Under the established existence of $\bm{B^{*}}$, the universal approximation capability of deep networks [cybenko1989approximation, hornik1989multilayer, leshno1993multilayer, kidger2020universal, yarotsky2022universal] and extensive validation experiments, our method is theoretically grounded and experimentally validated, providing an interpretable foundation and clarifying the theoretical upper bound. Our main contributions are as follows:
@@
  • Open Source. Code is publicly available on GitHub to foster community advancement.
@@ II-A Deep Learning-Based DoA Estimation
Figure 1: The overall structure of BeamformNet. The proposed learning paradigm is theoretically grounded and experimentally validated to approximate the optimal spatial filter via implicit spatial signal focusing and noise suppression, and the design of BeamformNet is based on beamforming principles.
@@ II-B Beamforming
Beamforming-based DoA estimation methods fall into static [krim1996two, capon2005high] and dynamic [zheng2023interpretable, deng2025spatial] categories: the former uses fixed closed-form expression filters, while the latter employs iterative optimization. They can also be classified as non-adaptive [krim1996two] or adaptive [capon2005high, deng2025spatial], depending on whether filter weights are adjusted based on received signals. In terms of implementation methodology, existing approaches can be divided into model-driven methods [krim1996two, capon2005high, deng2025spatial] and neural network-based methods [al2022review, zheng2023interpretable, hamza2024sparse]. In addition, beamforming algorithms can be divided into different methods based on different optimization objectives. CBF aims to maximize the Signal-to-Noise Ratio (SNR) under the assumption of white noise, prioritizing robustness but sacrificing resolution [krim1996two]. Adaptive methods like MVDR [capon2005high], on the other hand, seek to minimize the total output power subject to a distortionless constraint in the look direction. This objective implicitly forces the filter to place nulls in the directions of strong interference, thereby maximizing the Signal-to-Interference-plus-Noise Ratio (SINR). Extensions such as LCMV [1450747] introduce multiple linear constraints to shape the beam pattern further. Our method is a dynamic, adaptive beamforming approach, whose optimization objective is to find the optimal spatial filter, implemented via deep neural networks: its dynamic nature arises from iterative data training, and its adaptivity comes from generating spatial filters in response to input array signals.
II-C Exploiting Signal Sparsity
The spatial sparsity of signal sources—where signals occupy only a few angular directions among a fine grid—has inspired a class of Sparse Signal Recovery (SSR) methods [1614066, 1580791, 1468495, yang2018sparse, huang2025efficient]. These methods typically formulate DoA estimation as an inverse problem regularized by sparsity-inducing norms (e.g., $\ell_{1}$-norm or $\ell_{0}$ norm). Algorithms such as Orthogonal Matching Pursuit (OMP) [258082, 342465, 1337101] and Sparse Bayesian Learning (SBL) [2001Sparse] utilize this sparsity prior to achieve super-resolution performance, breaking the Rayleigh limit [van2002optimum]. While effective, traditional sparse reconstruction methods require iterative numerical optimization (e.g., using CVX [grant2008cvx] or ADMM [boyd2011distributed]), which is computationally expensive and difficult to deploy in real-time systems. In this work, we integrate the sparsity prior into the deep learning framework. By designing the network to output a spatial spectrum that matches the sparse distribution of ground-truth sources, BeamformNet effectively learns to perform sparse signal recovery in a forward-pass manner, combining the high resolution of sparsity-based methods with the inference speed of neural networks.
@@ III-C Beamforming and the Existence of the Optimal Spatial Filter
DoA is estimated from the spectral peak of the spatial energy spectrum $\bm{P}=[p_{1},...,p_{i},..,p_{R}]^{T}$, where $p_{i}=\bm{{b_{i}}^{H}XX^{H}b_{i}}$ with $p_{i}$ is the $i$-th element of $\bm{P}$ and $\bm{b_{i}}$ is the $i$-th column vector of matrix $\bm{B}^{H}$. For example, in the MVDR algorithm, it provides an analytical solution:
@@ IV-A Context Extraction in Time and Spatial Domains
To preserve the phase, BeamformNet splits $\bm{A}$ and $\bm{X}$ into real and imaginary parts, concatenates them, and feeds each into separate bidirectional RNNs to extract contextual features in the time and spatial domains in order to adaptively generate the spatial filter $\bm{B}$. Some works [zhou2018direction, su2024co, ma2024doa] show that virtual array elements can enhance array element dimensions for better DoA estimation. Accordingly, we expand the array element dimensions of $\bm{A}$ and $\bm{X}$ from $2M$ to $E$ ($E>2M$) after passing through the RNN, yielding an enriched representation that captures both temporal and spatial contexts with additional array element dimensions:
@@ IV-D Loss function
After obtaining the spatial energy spectrum $\bm{P}$, one possible approach is to directly learn according to Formula (12), which calculates the error between the obtained spatial energy spectrum $p_{i}=\bm{b_{i}^{H}XX^{H}b_{i}}$ and the sparse spatial signal energy spectrum $s_{i}(t)||^{2}$, such as using MSE (Mean Squared Error). Because our goal is to solve DoA, we are not interested in energy. So we directly solve for the direction in which the target angle is located. The DoAs obtained by searching the spatial energy spectrum $\bm{P}$ through spectral peaks cannot be directly used to compute gradients for backpropagation. In order to enable BeamformNet to backpropagate gradients for training, we apply the tanhtanh activation function to $\bm{P}$. Since each element in $\bm{P}$ lies in the range $0,+\infty)$, applying tanhtanh maps the signal energy in each grid cell to a probability distribution over $[0,1]$. Accordingly, BeamformNet formulates model training as a multi-label classification problem, which maximizes the probability of the grid where the target angle is located.
The above process uses the tanhtanh function (monotonically increasing) to convert $\bm{P}$ into probability and utilizes the sparsity of $\bm{S}$ for classification, which implies that the direction energy of the signal in $\bm{P}$ is higher (signal focusing), while other directions are lower or zero (noise suppression). It is equivalent to letting the network implicitly learn Formulas (13) and (14) (see Section VI-I for more details), although our network does not explicitly learn these two formulas. It is worth noting that although we do not directly and strictly constrain the network to adhere to Formula (12) (since we impose no constraints on energy), the suppression of noise and separation of signals compel BeamformNet to approximate Formula (12), thereby achieving near-optimal spatial filtering.
@@ VI-C Signal-to-Noise Ratio
The SNR measures signal strength relative to noise. A higher SNR indicates better signal quality and more accurate information transmission. We studied SNR’s impact on BeamformNet under $M=16$, $K=4$, $T=200$, and incoherent sources, comparing it with other methods. Figure 7 shows that DA-MUSIC performs best at low SNRs (-20 dB and -15 dB), while SubspaceNet and BeamformNet show comparable results. When the SNR exceeds -15 dB, BeamformNet outperforms all other methods, especially at high SNRs. We attribute BeamformNet’s weaker performance under extremely low SNRs to the fact that, although it suppresses noise at low SNRs (as shown in Table V), the noisy array received signal $\bm{X}$ impacts BeamformNet’s ability to separate and focus signals (Fomula 27). Enhancing BeamformNet’s signal separation and focusing performance under extremely low SNRs represents a promising area for future research. Nevertheless, overall, BeamformNet displays strong performance across the range of SNR conditions.
@@ VII Conclusion
This paper introduces BeamformNet, a novel deep learning-based beamforming method for DoA estimation. Drawing inspiration from traditional model-driven beamforming techniques, BeamformNet’s architecture and learning paradigm are theoretically founded on the approximation of an optimal spatial filter, enabling implicit signal focusing and noise suppression to achieve near-optimal DoA accuracy. Comprehensive experiments on simulated datasets and real-world near-field broadband speech localization benchmarks validate BeamformNet’s superior robustness and performance, surpassing state-of-the-art methods across diverse scenarios, including varying numbers of sources, snapshots, SNRs, angular separations, array elements, and mismatches—even under coherent sources and limited snapshots. These results underscore the efficacy of leveraging deep neural networks to approximate optimal solutions in array signal processing. Future research will aim to enhance the model’s capabilities in extreme conditions, such as single-snapshot and ultra-low-SNR environments.
@@
[Uncaptioned image] Kele Xu (Senior Member, IEEE) received the doctorate degree from Paris VI University, Paris, France, in 2017. He is currently an Associate Professor with the School of Computer Science, National University of Defense Technology, Changsha, China. His research interests include audio signal processing, machine learning, and intelligent software systems. He is the associate editor for IEEE Transactions on Circuits and Systems for Video Technology and Guest editor for Science Partner Journal Cyborg and Bionic Systems. He has (co-)authored more than 100 publications in peer reviewed journals and conference proceedings, including IEEE TASLP, TMI, ICML, CVPR, NeurIPS, ICLR, AAAI, IJCAI, SIGIR, ASE, ACM MM, ICASSP.